OpenStudy (anonymous):

Let f(x) = x^2 – 16. Find f^–1(x).

OpenStudy (anonymous):

anything to a negative power can be written with a positive power if it is moved to the denominator [x^(-2)]=1/(x^2) what do you think f^(-1)(x) is?

OpenStudy (anonymous):

is it 1/x^2-16?

OpenStudy (anonymous):

yes, 1/(x^2-16)

OpenStudy (dumbcow):

haha no that is not what they are looking for. \[f^{-1}(x) \] refers to the inverse function i hate the notation because it does look like an exponent, however f is not a variable its a function

OpenStudy (dumbcow):

basically you have to solve for x here. let f(x) be y \[y = x^{2} -16\] \[x = \pm \sqrt{y+16}\] \[\rightarrow f^{-1}(x) = \sqrt{x+16}\] important note, range of f(x) is y>-16 so domain of inverse must be x>-16